OpenStudy (anonymous):

Integrate e^[x] between limit 0 to 3

OpenStudy (amistre64):

there are no bad spots to worry about; so what does e^x come from? since it is a derivative of some function

OpenStudy (amistre64):

\[f(x)=\int^{b}_{a} f'(x)\ dx=f(b)-f(a)\]

OpenStudy (amistre64):

prolly some bad notation, but its off the cuff ;)

OpenStudy (anonymous):

this is int exp raise to power greatest integer x from 0 to 3

OpenStudy (lalaly):

do u know how to integrate e^x?

OpenStudy (anonymous):

this is only e^x .

OpenStudy (amistre64):

if i read you right, yes \[\int e^x = e^x\]

OpenStudy (amistre64):

\[e^3-e^0=\ ?\]

OpenStudy (amistre64):

greatest integer .... is that the ceiling function?

OpenStudy (amistre64):

if so, we might have to do this in pieces

OpenStudy (amistre64):

the wolf says greatest integer is the floor function ....

OpenStudy (amistre64):

the area can be considered as adding up the area of the rectangular spaces beneath the graph. using midpoints to make sure we are in each section gives me: 1*(e^[0.5]+e^[1.5]+e[2.5]) produces the same results

OpenStudy (zarkon):

\[\int_{0}^{3}e^{[x]}dx=\sum_{n=0}^{2}\int_{n}^{n+1}e^{[x]}dx\] \[=\sum_{n=0}^{2}\int_{n}^{n+1}e^{n}dx=\sum_{n=0}^{2}e^{n}(n+1-n)dx=\sum_{n=0}^{2}e^{n}\] \[=e^0+e^1+e^2\]